OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

is it (6,12)

OpenStudy (anonymous):

thats wat i mean insert 6 then get values

OpenStudy (anonymous):

as both equations are equal to y thus equating them we get 2x=x+6 2x-x=6 x=6 now put this value of x in any of these two equations y=2(6) y=12 thus solution is (6 , 12)